# On some practical challenges of conformal prediction

### v1

@@ 4 Illustration
Throughout this section, we will let ${\sf N}(\mu,\sigma^{2})$ denote the normal distribution with mean $\mu$ and variance $\sigma^{2}$, and let ${\sf Unif}(a,b)$ denote the uniform distribution supported on $(a,b)$.
Recall the standard linear regression model
$Y=X\beta+\sigma\varepsilon,$ (8)
where $Y$ is a $n$-dimensional vector of response variables, $X$, called data/design matrix, is an $n\times p$ matrix of observations of predictors, $\varepsilon$ is an $n$-dimensional vector of iid standard normal errors, $\beta$ is a $p$-dimensional vector of regression coefficients, and $\sigma>0$ is the scale parameter; if the model includes an intercept term, then the first column of $X$ consists of a $n$-vector of 1s and $X$ will be an $n\times(p+1)$ matrix. If we fit this model with the least squares estimation, the $100(1-\alpha)$% prediction interval for $Y_{n+1}$ at $X_{n+1}=x$, is
$x^{T}\hat{\beta}\pm t_{n-p}(\alpha/2)\,\hat{\sigma}\,\{1+x^{\top}(X^{\top}X)^{-1}x\}^{1/2},$ (9)
where $\hat{\beta}$ is the least squares estimator of $\beta$, $\hat{\sigma}$ is the residual standard error, $t_{\nu}(\alpha)$ denotes the $(1-\alpha)^{\text{th}}$ quantile of the (central) Student-t distribution with $\nu$ degrees of freedom. For $0<\alpha<1$, we will use $q_{\alpha}$ to denote the $(1-\alpha)^{\text{th}}$ quantile of $Y$.
@@
For $\alpha=0.1$, we generate $N=5,000$ random samples of size $n=1,001$ from the following model:
$Y=X_{1}+X_{2}+\epsilon,$
@@
where $X_{1},$ $X_{2}$, and $\epsilon$ are independent, and $X_{1}\sim{\sf N}(0,2)$, $X_{2}\sim{\sf N}(0,1)$, and $\epsilon\sim{\sf N}(0,\sqrt{0.2})$. For each sample, the response values of the first $1,000$ sample points and all the $1,001$ values of the two features are used to construct the conformal prediction intervals in the above three theorems as well as the linear model prediction intervals of the same three shapes. The two prediction intervals in Theorem 3 and Theorem 4 are independent of $\eta$. For the bounded prediction interval in Theorem 5, we take $\eta=\max_{1\leq i\leq n}\{2(\sqrt{\max\{0,c_{i}^{2}-d_{i}\}+1}-c_{i})\}$. Then, the 1,0011,001st response value is treated as the future response value we want to predict. We estimate the coverage probability of each prediction interval as $K/N$ where $K$ is the number of times it contains the 1,0011,001st response value. For the bounded (i.e., two-sided) $100(1-\alpha)\%$ prediction interval, we also calculate the ratio of its expected length to the length of the oracle interval $(q_{1-\alpha/2}-q_{\alpha/2})$. Table 1 summarizes the results.
Prediction interval form Linear Model Conformal Prediction
$\quad\quad(-\infty,a)$ 0.9038 0.9062
$\quad\quad(a,\infty)$ 0.9052 0.9034
$\quad\quad(a,b)$     0.9050 (0.1187)     0.9080 (1.1922)
Table 1: Coverage probabilities (ratio of interval length, if applicable) of the $90\%$ prediction intervals in Example A, based on the linear model and conformal prediction (Theorem 3, Theorem 4, and Theorem 5).
All three types of linear model prediction intervals and conformal prediction intervals achieve the nominal coverage probability. For conformal prediction intervals, this is no surprise because conformal prediction intervals are provably finite-sample valid. Since the linear model is well-specified in this example, the excellent performance of the linear model prediction intervals is expected. The bounded linear model prediction interval given by (9) is more efficient than the bound conformal prediction interval, though both are efficient. Since the conformal prediction interval is distribution-free, it is expected to be more conservative than the linear model prediction interval when the linear model is correct. Note that the oracle interval is based on information of $Y$ only, but the linear model prediction interval, given by (9), is constructed using information from both the response variable and predictors. Intuitively, when the variance of the noise (i.e, the error term $\epsilon$) is dominated by the variance of the predictors (Here $\mathsf{V}(X_{1}+X_{2})=3$ is much larger than $0.2=\mathsf{V}(\epsilon)$.), the oracle interval will be less efficient than the linear model prediction interval.
@@
We perform the same simulation with the same values of $\alpha$, $N$ and $n$ as in Example A except that data are generated from the following model:
$Y=X_{1}+X_{2}+\epsilon,$
where $X_{1},X_{2}$, and $\epsilon$ are independent and $X_{1}\sim{\sf N}(0,2)$, $X_{2}\sim{\sf N}(0,1)$, and $\epsilon\sim{\sf Unif}(-0.6,0.6)$. The results are summarized in Table 2.
Prediction interval form Linear Model Conformal Prediction
$\quad\quad(-\infty,a)$ 0.8750 0.9052
$\quad\quad(a,\infty)$ 0.8742 0.9034
$\quad\quad(a,b)$     0.9559 (0.1987)     0.9150 (1.1015)
Table 2: Coverage probabilities (ratior of interval length, if applicable) of the $90\%$ prediction intervals in Example B, based on the linear model and conformal prediction (Theorem 3, Theorem 4, and Theorem 5).
In this case, the linear model is incorrect. Therefore, the two one-sided linear model prediction intervals fail to provide adequate coverage. The bounded linear model prediction interval is generally not expected to provide adequate coverage when the linear model is wrong. It achieves the nominal coverage probability here simply because the probability mass of the distribution of the error term $\epsilon$ is concentrated in the interval (0.6,0.6)(-0.6,0.6). The three conformal prediction intervals all achieve the nominal coverage probability, as anticipated. While the bounded conformal prediction interval is a bit conservative, it is efficient when compared to the oracle prediction interval.
@@ 5 Concluding remarks
Conformal prediction is a powerful general strategy for creating finite-sample validity prediction intervals. For any non-conformity measure, the corresponding conformal prediction region is guaranteed to be finite-sample valid. However, the determination of a conformal prediction region generally requires a data scientist to evaluate the plausibility function for infinitely many values, which cannot be accomplished in practice. In the prior literature, many authors approximated the conformal prediction regions by evaluating the plausibility function for only a grid of possible values. The resulting region, not the same as the conformal prediction region, no longer has provable finite-sample validity, let alone two other challenges: (i) the computation required can still be prohibitively expensive, and (ii) the resulting prediction region might not be of a desired shape. While confronted with these practical challenges, some colleagues seem to believe that a monotonic nonconformity measure will resolve these issues. Our investigation showed that this is false, among other insights into the relationship between the monotonicity of the non-conformity measures, the monotonicity of the plausibility function, the shape of the conformal prediction region, and the exact determination of the conformal prediction regions. Our investigation also showed that it is challenging to give a hard-and-fast rule for choosing a non-conformity measure so that this issue can be avoided and the resulting conformal prediction region can be of a desired shape.
Driven by the principle of parsimony, we propose a non-conformity measure based on a multivariate polynomial of degree two. When we use the proposed non-conformity measure, we can not only avoid three common practical challenges in conformal prediction but also generate conformal prediction intervals of three common shapes. Of course, our proposal is by no means the only method a data scientist can employ to address the three common practical challenges. But it is simple and easy to implement. However, it does not work in higher-dimension cases. For example, if $Y\in\mathbb{R}^{p}$ for $p\geq 2$, our proposal may not apply, since there is no natural order on $\mathbb{R}^{p}$ for $p\geq 2$.
@@ 6 Appendix
This appendix documents the key results in the unsupervised learning.
@@ 6.1 Notation and setup
A similar strategy exists for the unsupervised learning setting. Suppose our observations consist of a sequence of exchangeable random variables $X_{1},X_{2},\ldots$, where each $X_{i}$ follows a distribution $\mathsf{P}$. We want to create a prediction interval of the next observation $X_{n+1}$, based on past observations of $X^{n}=\{X_{1},\ldots,X_{n}\}$. The conformal prediction algorithm in this case is the following Algorithm 2:
@@ 5   Compute $\mathsf{pl}_{x^{n}}(x)=(n+1)^{-1}\sum_{i=1}^{n+1}1\{\mu_{i}\geq\mu_{n+1}\}$;
8Return $\mathsf{pl}_{z^{n}}(x)$ for each $x$;
@@ A $100(1-\alpha)\%$ conformal prediction region can be constructed as follows:
$C_{\alpha}(X^{n})=\{x:\mathsf{pl}_{Z^{n}}(x)>\alpha\},$ (10)
where $0<\alpha<1$. The finite-sample validity still holds:
$\mathsf{P}^{n+1}\{X_{n+1}\in C_{\alpha}(X_{n+1})\}\geq 1-\alpha\quad\text{for all $(n,\mathsf{P})$},$
@@
Let $B=\{x_{1},\ldots,x_{n}\}$ be a bag of observations of size $n$. For $i=1,\ldots,n$, let $(x_{1}\ldots,x_{n})$ be the observations in $B$. Suppose $x$ is a provisional value of a future observation of $X$ to be predicted. In this case, a non-conformity measure $M$ is said to be monotonically increasing if
@@
In the unsupervised learning case, (3) corresponds to
@@
  7. (VII’)
    Is the monotonicity of $M$ a necessary condition for the exact determination of $C_{\alpha}(Z^{n})$?
@@
    Is (11) a necessary condition for exact determination of $C_{\alpha}(Z^{n})$? (Note that the converse is true: (11) implies $C_{\alpha}(Z^{n})$ is a one-sided interval; hence it implies the exact determination of $C_{\alpha}(Z^{n})$.)
@@
    Is monotonicity of the plausibility function $\mathsf{pl}_{z^{n}}$ a necessary condition for the exact determination of $C_{\alpha}(Z^{n})$? (Note that the converse holds, i.e., monotonicity of the plausibility function implies $C_{\alpha}(X^{n})$ is a one-sided interval; hence, it implies the exact determination of $C_{\alpha}(X^{n})$.)
6.2 Answers to Questions (I’)—(IX’)
Answers to Questions (I’)—(IX’) follow from answers to Questions (I)—(IX) since we can simply ignore the predictors and treat the response variable as the random variable of interest in unsupervised learning. For the sake of completeness, we give the details below.
@@ 6.2 Answers to Questions (I’)—(IX’)
Example 10.
Let $M(B,x)=\min\{x_{1},\ldots,x_{n}\}+x$. Then $M$ is monotonic. Also,
$\mu_{i}=M(X^{n+1}\backslash X_{i},X_{i})=\min\{X_{1},\ldots,X_{i-1},X_{i+1},\ldots,X_{n},X_{n+1}\}+X_{i},\quad i=1,\ldots,n,n+1.$
In particular, $\mu_{n+1}=\min\{X_{1},\ldots,X_{n}\}+X_{n+1}$. Thus,
$\mu_{i}\geq\mu_{n+1}$ $\Longleftrightarrow$ $\min\{X_{1},\ldots,X_{i-1},X_{i+1},\ldots,X_{n},X_{n+1}\}+X_{i}$ (12)
$\geq\min\{X_{1},\ldots,X_{n}\}+X_{n+1}$
$\Longleftrightarrow$ $\min\{m_{i},X_{n+1}\}+X_{i}\geq\min\{m_{i},X_{i}\}+X_{n+1},$
where $m_{i}=\min\{X_{1},\ldots,X_{i-1},X_{i+1},\ldots,X_{n}\}$. If $m_{i}\geq X_{n+1}$, the last inequiality of (12) becomes
$X_{n+1}+X_{i}\geq\min\{m_{i},X_{i}\}+X_{n+1},$
which is obviously true. When $m_{i}<X_{n+1}$, (12) is equivalent to
$m_{i}+X_{i}\geq\min\{m_{i},X_{i}\}+X_{n+1},$
i. e., $m_{i}<X_{n+1}\leq X_{i}+m_{i}-\min\{m_{i},X_{i}\}$. If $m_{i}\leq X_{i}$, this becomes $m_{i}<X_{n+1}\leq X_{i}$. We cannot have $m_{i}>X_{i}$ because it would imply $m_{i}<X_{n+1}\leq m_{i}$, which is absurd. Therefore, (11) does not hold in this case.
@@
Example 11.
Let $M(B,x)=(x_{1}^{2}+\ldots+x_{n}^{2})+x^{2}+x$. Clearly, $M$ is not monotonic. We have
$\mu_{i}=M(X^{n+1}\backslash\{X_{i}\},X_{i})=\sum_{j=1,j\neq i}^{n+1}X_{j}^{2}+X_{i}^{2}+X_{i},\quad j=1,\ldots,n,n+1.$
Hence, $\mu_{i}\geq\mu_{n+1}$ if and only if
$\sum_{j=1,j\neq i}^{n+1}X_{j}^{2}+X_{i}^{2}+X_{i}\geq\sum_{j=1}^{n}X_{j}^{2}+X_{n+1}^{2}+X_{n+1},$
which is equivalent to $X_{n+1}\leq X_{i}$.
@@
Example 12.
Consider Example 11. Here $M$ is not monotonic. However, (11) holds, which implies $C_{\alpha}(X^{n})=(-\infty,X_{(k)})$.
@@
Example 13.
Let $M(B,x)=x_{1}^{2}+\ldots+x_{n}^{2}+x$. Then $M$ is monotonic. We have
$\mu_{i}=M(X^{n+1}\backslash\{X_{i}\},X_{i})=\sum_{j=1,j\neq i}^{n+1}X_{j}^{2}+X_{i}\quad i=1,\ldots,n,n+1.$
Thus,
$\mu_{i}\geq\mu_{n+1}\Longleftrightarrow\sum_{j=1,j\neq i}^{n+1}X_{j}^{2}+X_{i}\geq\sum_{j=1}^{n}X_{j}^{2}+X_{n+1}.$
It follows that
$\mu_{i}\geq\mu_{n+1}\Longleftrightarrow X_{n+1}^{2}+X_{i}\geq X_{i}^{2}+X_{n+1},$
implying $X_{n+1}\in(-\infty,\min\{X_{i},1-X_{i}\})\cup(\max\{X_{i},1-X_{i}\},\infty)$.
Answer to Question (V’): No. (11) is not a necessary condition for $C_{\alpha}(X^{n})$ to be an interval.
Example 14.
Consider Example 10. We know that (11) does not hold in this case. Now note that (12) implies that
$C_{\alpha}(X^{n})=(-\infty,a_{(k)}),$
where $a_{i}=X_{i}+\min\{X_{1},\ldots,X_{i-1},X_{i+1},\ldots,X_{n}\}-\min\{X_{1},\ldots,X_{n}\}$. Therefore, $C_{\alpha}(X^{n})$ is a one-sided interval.
Answer to Question (VI’): Yes. Monotonicity of the plausibility function $\mathsf{pl}_{x^{n}}$ a necessary condition for $C_{\alpha}(X^{n})$ to be a one-sided interval.
Proof.
Suppose $C_{\alpha}(X^{n})$ is a one-sided interval. Without loss of generality, we assume the plausibility function $\mathsf{pl}_{x^{n}}$ is not monotonically increasing. Then, there are three numbers $a<b<c$ such that $\mathsf{pl}_{x^{n}}(b)>\mathsf{pl}_{x^{n}}(a)$ and $\mathsf{pl}_{x^{n}}(b)>\mathsf{pl}_{x^{n}}(c)$. Thus, for any confidence level $\alpha$ such that $\max\{\mathsf{pl}_{x^{n}}(a),\mathsf{pl}_{x^{n}}(c)\}<\alpha<\mathsf{pl}_{x^{n}}(b)$, we will have $b\in C_{\alpha}(X^{n})$ but $a\not\in C_{\alpha}(X^{n})$ and $c\not\in C_{\alpha}(X^{n})$. Therefore, $C_{\alpha}(X^{n})$ cannot be a one-sided interval. ∎
Answer to Question (VII’): No. Monotonicity of $M$ is not necessary condition for exact determination of $C_{\alpha}(X^{n})$.
Example 15.
Consider Example 11. Here $C_{\alpha}(X^{n})=(-\infty,X_{(k)})$. Hence, we can determine $C_{\alpha}(Z^{n})$ exactly, though $M$ is not monotonic.
@@
Example 16.
Consider Example 14.
Answer to Question (IX’): No. Monotonicity of the plausibility function $\mathsf{pl}_{z^{n}}$ is not a necessary condition for exact determination of $C_{\alpha}(X^{n})$.
Example 17.
Consider Example 13.
6.3 Proposed strategy
@@
If $\lfloor(n+1)\alpha\rfloor\leq 1$, then the $(1-\alpha)\%$ prediction region $C_{\alpha}(X^{n})$ given by (10) is $\mathbb{R}$.
@@
where $B=\{x_{1},\ldots,x_{n}\}$ and $\lambda,\theta$, and $\kappa$ are constants to be decided by the user. Following the same line of reasoning as in the previous section, we can see that the following three theorems hold. Note that Theorems 7 and 8 recover the traditional non-parametric one-sided prediction intervals based on order statistics (e.g., Wiks 941; Fligner and Wolfe 1976; Frey 2013).
@@
Suppose $0<\alpha<1$, $\lfloor(n+1)\alpha\rfloor\geq 2$, and the non-conformity measure is given by (13). If $\lambda=1$, $\theta=0$, $\kappa\neq 0$, then the $100(1-\alpha)\%$ conformal prediction region $C_{\alpha}(X_{n})$ is the bounded interval $(a_{(r_{1})},b_{(r_{2})})$, where $a_{i}=\min\{-(\kappa+X_{i}),X_{i}\}$ and $b_{i}=\max\{-(\kappa+X_{i}),X_{i}\}$ for $1\leq i\leq n$.

### v2

@@ 4 Illustration
Throughout this section, we will use the following notation:
  1. $\bullet$
    ${\sf Binom}(k,p)$ denotes the binomial distribution with trial number parameter $k$ and success probability $0<p<1$;
  2. $\bullet$
    ${\sf Gamma}(a,b)$ represents the gamma distribution with shape parameter $a$ and rate parameter $b$, i.e., the gamma distribution whose probability density function is
    $f(x)=\frac{b^{a}}{\Gamma(a)}x^{a-1}e^{-bx},\quad x>0,$
    where $\Gamma(\alpha)=\int_{0}^{\infty}t^{\alpha-1}e^{-t}dt$ is the Gamma function;
  3. $\bullet$
    ${\sf N}(\mu,\sigma^{2})$ stands for the normal distribution with mean $\mu$ and variance $\sigma^{2}$;
  4. $\bullet$
    ${\sf Pareto}(\eta,\beta)$ symbolizes the Type-II Pareto distribution with the probability density function
    $g(x)=\frac{\eta\beta^{\eta}}{(x+\beta)^{\eta+1}},\quad x>0.$
@@ 4. $\bullet$
For $\alpha=0.1$, we generate $N=3,000$ random samples of size $n=101$ from the following model:
$Y=X_{1}+X_{2}+\varepsilon,$
@@
where $X_{1},$ $X_{2}$, and $\varepsilon$ are independent, and $X_{1}\sim{\sf N}(0,2)$, $X_{2}\sim{\sf N}(0,1)$, and $\varepsilon\sim{\sf N}(0,0.4)$. For each sample, the response values of the first $100$ sample points and all the $101$ values of the two features are used to construct the conformal prediction intervals in the above three theorems. We let $(\delta+\gamma)/1-\eta=1$ and 1-1 in Theorem 3 and Theorem 4, respectively. For the bounded prediction interval in Theorem 5, we take $\delta=1+0.8\underset{i^{\prime\prime}}{\min}\left\{Y_{i^{\prime\prime}}^{2}/\sum_{j=1}^{p}(X_{(n+1)j}-X_{i^{\prime\prime}j})\right\}$, where $\sum_{j=1}^{p}(X_{(n+1)j}-X_{i^{\prime\prime}j})<0$. Then, the 101101st response value is treated as the future response value we want to predict. We estimate the coverage probability of each prediction interval as $K/N$ where $K$ is the number of times it contains the 101101st response value. For the $(a,b)$-shaped $100(1-\alpha)\%$ prediction interval, we report its mean length as a measure of its efficiency. For the $(-\infty,a)$-shaped prediction interval, we measure its efficiency using $\mathsf{E}[a]$. Likewise, the efficiency of the $(a,\infty)$-shaped prediction interval is measured with $\mathsf{E}[a]$. We do the same for split conformal prediction with least squares absolute test residuals/predictive errors in Lei et al. (2018) and the transformation method with the linear function $h(t_{1},t_{2})=1+t_{1}+t_{2}$ in Hong (2016b). That is, we let $R_{i}$ be the $i$-th least squares absolute test predictive error in the split conformal algorithm in Lei et al. (2018) to obtain an equal-tailed bounded prediction interval. We obtain the two one-sided prediction intervals in the obvious way by using the same split conformal prediction algorithm with $R_{i}$ being the $i$-th test least squares predictive error. For the transformation method in Hong (2026b), we shift $(-\infty,W_{(r_{1})})$, $(W_{(r_{2})},+\infty)$ and $(W_{(r_{3})},W_{(n-r_{3}+1)})$ by $h(X_{n+1,1},X_{n+1,2})$ to obtain the three prediction intervals for $Y_{n+1}$. Table 1 summarizes the results.
Prediction Interval Shape Full Split Transformation
$\quad\quad(-\infty,a)$ 0.91 (0.82) 0.91 (0.81) 0.91 (0.82)
$\quad\quad(a,\infty)$ 0.91 (-0.86) 0.90 (-0.91) 0.91 (-0.88)
$\quad\quad(a,b)$ 0.92 (6.16) 0.90 (2.18) 0.90 (2.14)
Table 1: Coverage probabilities (efficiency) of three types of the $90\%$ prediction intervals in Example A, based on full conformal prediction with proposed non-conformity measures, split conformal prediction with least square absolute test predictive errors, and the transformation method with a linear transformation.
For all three desired shapes, prediction intervals based on the proposed method, split conformal prediction, and the transformation method all achieve the nominal cover probability. This is no surprise because these prediction intervals are all provably finite-sample valid. For the $(-\infty,a)$-shaped prediction interval, the three methods are almost equally efficient. The proposed method slightly outperforms the other two methods for the $(a,\infty)$-shaped prediction interval. However, for the $(a,b)$-shape prediction intervals, the proposed method is much more conservative than the other two methods. Our explanation for this is that when the true data-generating model is a bona fide linear model, the information needed for constructing a finite-sample valid prediction interval is mainly contained in the order of data. Thus, the non-conformity measure in Theorem 5, a quadratic functional of data, might not efficiently utilize the information in the data to create a finite-sample valid prediction interval.
@@
We perform the same simulation with the same values of $\alpha$, $N$ and $n$ as in Example A, except that (a) data are generated from the following model:
$Y=X_{1}^{2}-X_{2}^{3}+X_{3}+\varepsilon,$
where $X_{1},X_{2}$, and $\varepsilon$ are independent, and $X_{1}\sim{\sf N}(0,4)$, $X_{2}\sim{\sf Gamma}(2,1)$, $X_{3}\sim{\sf Binom}(2,0.4)$, and $\varepsilon\sim{\sf Pareto}(2,7)$, and (b) we use $h(t_{1},t_{2},t_{3})=1+t_{1}+t_{2}+t_{3}$ in the transformation method. The results are summarized in Table 2.
Prediction Interval Shape Full Split Transformation
$\quad\quad(-\infty,a)$ 0.90 (9.14) 0.91 (10.11) 0.90 (9.14)
$\quad\quad(a,\infty)$ 0.90 (-61.19) 0.90 (-48.21) 0.91 (-68.12)
$\quad\quad(a,b)$ 0.90 (118.41) 0.91 (99.30) 0.90 (132.29)
Table 2: Coverage probabilities (efficiency) of three types of the $90\%$ prediction intervals in Example B, based on full conformal prediction with proposed non-conformity measures, split conformal prediction with least square absolute test predictive errors, and the transformation method with a linear transformation.
This time the true data-generating model is not a classical linear model, since the response variable $Y$ is a cubic polynomial of the three predictors and $\mathsf{E}[\varepsilon]\neq 0$. (Note that none of the three methods needs the “convenient” assumption $\mathsf{E}[\varepsilon]=0$.) Similar to what we have seen in Example A, all prediction intervals attain the nominal coverage probability. For the $(-\infty,a)$-shaped prediction intervals, the proposed method and the transformation method are tied, and they both outperform split conformal prediction. For the $(a,\infty)$-shaped prediction intervals as well as the $(a,b)$-shaped prediction intervals, split conformal prediction is more efficient than the other two methods, with the proposed method being slightly more efficient than the transformation method.
Example C
We repeat the same simulation with the same values of $\alpha$, $N$ and $n$ as in Example B, except that the true data-generating model is
$Y=X_{1}-X_{2}+e^{X_{3}}-\varepsilon,$
where $X_{1},X_{2}$, and $\varepsilon$ are independent, and $X_{1}\sim{\sf Pareto}(3,2)$, $X_{2}\sim{\sf N}(0,1)$, $X_{3}\sim{\sf Gamma}(2,2)$, and $\varepsilon\sim{\sf Binom}(10,0.4)$. The results are summarized in Table 2.
Prediction Interval Shape Full Split Transformation
$\quad\quad(-\infty,a)$ 0.90 (7.22) 0.90 (7.47) 0.90 (7.31)
$\quad\quad(a,\infty)$ 0.90 (-7.59) 0.90 (-7.69) 0.92 (-7.83)
$\quad\quad(a,b)$ 0.91 (18.17) 0.91 (21.78) 0.91 (19.84)
Table 3: Coverage probabilities (efficiency) of three types of the $90\%$ prediction intervals in Example C, based on full conformal prediction with proposed non-conformity measures, split conformal prediction with least square absolute test predictive errors, and the transformation method with a linear transformation.
All prediction intervals in Table 3 achieve their nominal coverage probability, as anticipated. In terms of efficiency, the proposed method outperforms split conformal prediction and the transformation method for all three interval shapes. This time split conformal prediction is the least efficient one among the $(a,b)$-shaped prediction interval.
If a data scientist only cares about finite-sample validity, then any of these three methods will do a good job. When it comes to efficiency, we are not prepared to assign supremacy to any one of them over the other two. However, the above three examples seem to suggest that split conformal prediction is likely to be an excellent choice when the response variable is truly linear in predictors or a polynomial of predictors, and that the proposed method and the transformation can work well when the true data-generating mechanism is complicated.
@@ 5 Concluding remarks
Within the full conformal prediction framework, a conformal prediction region is guaranteed to be finite-sample valid. However, the determination of a conformal prediction region generally requires a data scientist to evaluate the plausibility function at infinitely many values, which cannot be completed in practice. If we evaluate the plausibility function for only a grid of possible values, the resulting region, an approximation of the conformal prediction region, is no longer guaranteed to be finite-sample valid, let alone there are two other challenges: (i) the computation required can still be prohibitively expensive, and (ii) the resulting prediction region might not be of a desired shape. Some colleagues tend to believe that a monotonic nonconformity measure will overcome these challenges. Our investigation showed that this is false, among other insights into the relationship between the monotonicity of the non-conformity measures, the monotonicity of the plausibility function, the shape of the conformal prediction region, and the exact determination of the conformal prediction regions. Our investigation also showed that it is challenging to give a hard-and-fast rule for choosing a non-conformity measure so that this issue can be avoided and the resulting conformal prediction region can be of a desired shape.
Driven by the principle of parsimony, we propose a non-conformity measure based on a multivariate quadratic polynomial. With the proposed non-conformity measure, we can not only avoid three common practical challenges in conformal prediction but also generate conformal prediction intervals of three common shapes. Simple and easy to implement, our proposal enables a data scientist to both determine the prediction interval exactly and control its shape in the full conformal prediction framework.
Acknowledgments
We thank the two anonymous reviewers for many helpful comments and suggestions.
Conflict of interest
The authors declare no conflict of interest.
@@ 6 Appendix
This appendix briefly documents the key results in unsupervised learning.
@@ 6.1 Notation and setup
Suppose our observations consist of a sequence of exchangeable random variables $X_{1},X_{2},\ldots$, where each $X_{i}$ follows a distribution $\mathsf{P}$. We want to construct a prediction interval of the next observation $X_{n+1}$, based on past observations of $X^{n}=\{X_{1},\ldots,X_{n}\}$. The conformal prediction algorithm in this case is the following Algorithm 2:
@@ 5   Compute $\mathsf{pl}_{x^{n}}(x)=(n+1)^{-1}\sum_{i=1}^{n+1}1\{\mu_{i}\geq\mu_{n+1}\}$;
8Return $\mathsf{pl}_{x^{n}}(x)$ for each $x$;
@@ A $100(1-\alpha)\%$ conformal prediction region can be constructed as follows:
$C_{\alpha}(X^{n})=\{x:\mathsf{pl}_{X^{n}}(x)>t_{n}(\alpha)\},$ (10)
where $0<\alpha<1$ and $t_{n}(\alpha)=(n+1)^{-1}\lfloor(n+1)(1-\alpha)\rfloor$. The finite-sample validity still holds:
$\mathsf{P}^{n+1}\{X_{n+1}\in C_{\alpha}(X^{n})\}\geq 1-\alpha\quad\text{for all $(n,\mathsf{P})$},$
@@
Let $B=\{x_{1},\ldots,x_{n}\}$ be a bag of observations of size $n$. For $i=1,\ldots,n$, let $(x_{1}\ldots,x_{n})$ be the observations in $B$. Suppose $x$ is a provisional value of a future observation of $X_{n+1}$ to be predicted. In this case, a non-conformity measure $M$ is said to be monotonically increasing if
@@
In unsupervised learning, (4) corresponds to
@@
    Is the monotonicity of $M$ a necessary condition for the exact determination of $C_{\alpha}(X^{n})$?
  6. (VI’)
@@
    Is (11) a necessary condition for exact determination of $C_{\alpha}(X^{n})$? (Note that the converse is true: (11) implies $C_{\alpha}(X^{n})$ is a one-sided interval; hence it implies the exact determination of $C_{\alpha}(X^{n})$.)
@@
    Is monotonicity of the plausibility function $\mathsf{pl}_{x^{n}}$ a necessary condition for the exact determination of $C_{\alpha}(X^{n})$? (Note that the converse holds, i.e., monotonicity of the plausibility function implies $C_{\alpha}(X^{n})$ is a one-sided interval; hence, it implies the exact determination of $C_{\alpha}(X^{n})$.)
Answers to Questions (I’)—(IX’) follow from answers to Questions (I)—(IX) by ignoring the predictors and treating the response variable as the random variable of interest in unsupervised learning. For example, if we take $M(B,x)=\min\{x_{1},\ldots,x_{n}\}+x$, we will obtain an example to show the answer to Question (I’) is negative. For the sake of brevity, we give answers to Questions (I’)—(IX’) without proofs.
@@
Answer to Question (V’): No. Monotonicity of $M$ is not necessary condition for exact determination of $C_{\alpha}(X^{n})$.
Answer to Question (VI’): No. (11) is not a necessary condition for $C_{\alpha}(X^{n})$ to be an interval.
Answer to Question (VII’): Yes. Monotonicity of the plausibility function $\mathsf{pl}_{x^{n}}$ a necessary condition for $C_{\alpha}(X^{n})$ to be a one-sided interval.
@@
Answer to Question (IX’): No. Monotonicity of the plausibility function $\mathsf{pl}_{x^{n}}$ is not a necessary condition for exact determination of $C_{\alpha}(X^{n})$.
@@
If $\lfloor(n+1)\alpha\rfloor<1$, then the $(1-\alpha)\%$ prediction region $C_{\alpha}(X^{n})$ given by (10) is $\mathbb{R}$.
@@
where $B=\{x_{1},\ldots,x_{n}\}$ and $\lambda,\theta$, and $\kappa$ are constants to be decided by the user. Following the same line of reasoning as in the previous section, we can see that the following three theorems hold. Note that Theorems 8 and 9 recover the traditional non-parametric one-sided prediction intervals based on order statistics (e.g., Wiks 941; Fligner and Wolfe 1976; Frey 2013).
@@
Suppose $0<\alpha<1$, $\lfloor(n+1)\alpha\rfloor\geq 1$, and the non-conformity measure is given by (12). If $\lambda=1$, $\theta=0$, $\kappa\neq 0$, then the $100(1-\alpha)\%$ conformal prediction region $C_{\alpha}(X_{n})$ is the bounded interval $(a_{(r_{3})},b_{(n-r_{3}+1)})$, where $a_{i}=\min\{-(\kappa+X_{i}),X_{i}\}$ and $b_{i}=\max\{-(\kappa+X_{i}),X_{i}\}$ for $1\leq i\leq n$.